# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

### v2

@@ 5 HoTPP Benchmark
Method Mean length Next-item Long-horizon Acc (%) mAP (%) MAE OTD Val / Test T-mAP Val / Test (%) Transactions MostPopular 7.5 32.78$\pm$ 0.00 0.86$\pm$ 0.00 0.752$\pm$ 0.000 7.37$\pm$ 0.00 / 7.38$\pm$ 0.00 1.06$\pm$ 0.00 / 0.99$\pm$ 0.00 Last 5 4.7 19.60$\pm$ 0.00 0.87$\pm$ 0.00 0.924$\pm$ 0.000 7.40$\pm$ 0.00 / 7.44$\pm$ 0.00 2.46$\pm$ 0.00 / 2.49$\pm$ 0.00 IFTPP 11.0 34.08$\pm$ 0.04 3.47$\pm$ 0.01 0.693$\pm$ 0.000 6.88$\pm$ 0.01 / 6.90$\pm$ 0.01 5.82$\pm$ 0.27 / 5.88$\pm$ 0.13 IFTPP-K 10.1 33.69$\pm$ 0.03 3.25$\pm$ 0.01 0.698$\pm$ 0.001 7.18$\pm$ 0.00 / 7.19$\pm$ 0.00 4.42$\pm$ 0.15 / 4.43$\pm$ 0.16 RMTPP 7.5 34.15$\pm$ 0.07 3.47$\pm$ 0.02 0.749$\pm$ 0.005 6.86$\pm$ 0.01 / 6.88$\pm$ 0.01 7.08$\pm$ 0.16 / 6.69$\pm$ 0.12 RMTPP-K 7.3 33.63$\pm$ 0.07 3.24$\pm$ 0.02 0.749$\pm$ 0.001 7.10$\pm$ 0.00 / 7.11$\pm$ 0.00 5.82$\pm$ 0.05 / 5.52$\pm$ 0.13 NHP 9.4 35.43$\pm$ 0.07 3.41$\pm$ 0.01 0.696$\pm$ 0.002 6.97$\pm$ 0.01 / 6.98$\pm$ 0.01 5.59$\pm$ 0.07 / 5.61$\pm$ 0.05 AttNHP 7.6 31.12$\pm$ x.xx 1.21$\pm$ x.xx 0.717$\pm$ x.xxx 7.52$\pm$ x.xx / 7.50$\pm$ x.xx 1.71$\pm$ x.xx / 1.48$\pm$ x.xx ODE 9.1 35.60$\pm$ 0.06 3.34$\pm$ 0.06 0.695$\pm$ 0.002 6.96$\pm$ 0.01 / 6.97$\pm$ 0.01 5.53$\pm$ 0.08 / 5.52$\pm$ 0.13 HYPRO 6.9 34.26$\pm$ x.xx 3.46$\pm$ x.xx 0.758$\pm$ x.xxx 7.04$\pm$ x.xx / 7.05$\pm$ x.xx 7.79$\pm$ x.xx / 7.05$\pm$ x.xx MIMIC-IV MostPopular 10.3 4.77$\pm$ 0.00 2.75$\pm$ 0.00 14.52$\pm$ 0.00 19.82$\pm$ 0.00 / 19.82$\pm$ 0.00 0.55$\pm$ 0.00 / 0.54$\pm$ 0.00 Last 5 4.0 1.02$\pm$ 0.00 2.63$\pm$ 0.00 5.34$\pm$ 0.00 19.75$\pm$ 0.00 / 19.73$\pm$ 0.00 2.41$\pm$ 0.00 / 2.49$\pm$ 0.00 IFTPP 12.7 58.59$\pm$ 0.02 47.32$\pm$ 1.39 3.00$\pm$ 0.01 11.51$\pm$ 0.01 / 11.53$\pm$ 0.01 21.93$\pm$ 0.21 / 21.67$\pm$ 0.21 IFTPP-K 15.1 57.91$\pm$ 0.13 44.60$\pm$ 0.14 3.07$\pm$ 0.02 13.17$\pm$ 0.05 / 13.18$\pm$ 0.05 22.46$\pm$ 0.03 / 22.30$\pm$ 0.03 RMTPP 10.0 58.33$\pm$ 0.02 46.24$\pm$ 1.18 3.89$\pm$ 0.03 13.64$\pm$ 0.03 / 13.71$\pm$ 0.03 21.49$\pm$ 0.31 / 21.08$\pm$ 0.29 RMTPP-K 8.4 57.48$\pm$ 0.06 43.47$\pm$ 0.11 3.62$\pm$ 0.02 14.68$\pm$ 0.01 / 14.72$\pm$ 0.02 20.70$\pm$ 0.16 / 20.39$\pm$ 0.14 NHP 6.1 24.97$\pm$ 0.94 11.12$\pm$ 0.49 6.53$\pm$ 0.77 18.59$\pm$ 0.18 / 18.60$\pm$ 0.19 7.26$\pm$ 1.35 / 7.32$\pm$ 1.33 AttNHP 5.1 42.75$\pm$ 0.52 28.55$\pm$ 1.06 3.08$\pm$ 0.04 14.66$\pm$ 0.08 / 14.68$\pm$ 0.08 22.64$\pm$ 0.41 / 22.46$\pm$ 0.40 ODE 12.5 43.21$\pm$ 2.30 25.34$\pm$ 3.05 2.93$\pm$ 0.03 14.71$\pm$ 0.34 / 14.74$\pm$ 0.34 15.41$\pm$ 0.21 / 15.18$\pm$ 0.15 HYPRO 7.4 58.35$\pm$ x.xx 45.45$\pm$ x.xx 3.95$\pm$ x.xx 14.82$\pm$ x.xx / 14.87$\pm$ x.xx 16.94$\pm$ x.xx / 16.77$\pm$ x.xx Retweet MostPopular 28.0 58.50$\pm$ 0.00 39.85$\pm$ 0.00 18.82$\pm$ 0.00 174.9$\pm$ 0.0 / 173.5$\pm$ 0.0 25.15$\pm$ 0.00 / 23.91$\pm$ 0.00 Last 10 9.8 50.29$\pm$ 0.00 35.73$\pm$ 0.00 21.87$\pm$ 0.00 152.3$\pm$ 0.0 / 150.3$\pm$ 0.0 29.12$\pm$ 0.00 / 29.24$\pm$ 0.00 IFTPP 26.1 59.95$\pm$ 0.03 46.53$\pm$ 0.04 18.27$\pm$ 0.03 173.3$\pm$ 4.1 / 172.7$\pm$ 4.4 34.90$\pm$ 4.63 / 31.75$\pm$ 4.44 IFTPP-K 14.8 59.55$\pm$ 0.26 45.09$\pm$ 0.62 18.21$\pm$ 0.42 168.6$\pm$ 2.4 / 167.9$\pm$ 2.8 37.11$\pm$ 4.91 / 34.73$\pm$ 5.11 RMTPP 19.0 60.07$\pm$ 0.10 46.81$\pm$ 0.06 18.45$\pm$ 0.16 167.6$\pm$ 3.4 / 166.7$\pm$ 3.3 47.86$\pm$ 0.91 / 44.74$\pm$ 0.94 RMTPP-K 13.7 59.99$\pm$ 0.05 46.34$\pm$ 0.09 18.33$\pm$ 0.13 164.7$\pm$ 2.8 / 163.9$\pm$ 2.8 49.07$\pm$ 1.16 / 46.16$\pm$ 1.32 NHP 18.1 60.08$\pm$ 0.04 46.83$\pm$ 0.11 18.42$\pm$ 0.11 167.0$\pm$ 1.6 / 165.8$\pm$ 1.6 48.31$\pm$ 0.67 / 45.07$\pm$ 0.34 AttNHP 29.5 60.03$\pm$ 0.03 46.74$\pm$ 0.01 18.39$\pm$ 0.05 173.3$\pm$ 1.0 / 171.6$\pm$ 1.0 28.32$\pm$ 1.24 / 25.85$\pm$ 1.08 ODE 19.0 59.95$\pm$ 0.08 46.65$\pm$ 0.04 18.38$\pm$ 0.04 166.5$\pm$ 0.5 / 165.3$\pm$ 0.5 48.70$\pm$ 0.93 / 44.81$\pm$ 0.69 HYPRO 15.0 59.87$\pm$ x.xx 46.69$\pm$ x.xx 18.75$\pm$ x.xx 171.4$\pm$ x.x / 170.7$\pm$ x.x 49.90$\pm$ x.xx / 46.99$\pm$ x.xx Amazon MostPopular 20.6 33.46$\pm$ 0.00 9.58$\pm$ 0.00 0.304$\pm$ 0.000 7.20$\pm$ 0.00 / 7.18$\pm$ 0.00 8.59$\pm$ 0.00 / 8.31$\pm$ 0.00 Last 5 5.0 24.23$\pm$ 0.00 8.15$\pm$ 0.00 0.321$\pm$ 0.000 6.43$\pm$ 0.00 / 6.41$\pm$ 0.00 9.73$\pm$ 0.00 / 9.21$\pm$ 0.00 IFTPP 13.7 35.73$\pm$ 0.12 17.14$\pm$ 0.10 0.242$\pm$ 0.003 6.58$\pm$ 0.06 / 6.52$\pm$ 0.05 21.94$\pm$ 0.46 / 22.56$\pm$ 0.52 IFTPP-K 13.3 35.11$\pm$ 0.08 16.48$\pm$ 0.03 0.246$\pm$ 0.000 6.72$\pm$ 0.01 / 6.68$\pm$ 0.01 22.06$\pm$ 0.05 / 22.57$\pm$ 0.07 RMTPP 16.7 35.76$\pm$ 0.06 17.21$\pm$ 0.02 0.294$\pm$ 0.001 6.62$\pm$ 0.02 / 6.57$\pm$ 0.03 19.70$\pm$ 0.31 / 20.06$\pm$ 0.33 RMTPP-K 15.8 35.06$\pm$ 0.14 16.37$\pm$ 0.20 0.300$\pm$ 0.005 6.92$\pm$ 0.01 / 6.87$\pm$ 0.02 17.85$\pm$ 0.29 / 18.12$\pm$ 0.30 NHP 9.9 11.06$\pm$ 3.43 11.22$\pm$ 0.19 0.449$\pm$ 0.014 9.04$\pm$ 0.31 / 9.02$\pm$ 0.35 26.24$\pm$ 0.36 / 26.29$\pm$ 0.55 AttNHP 18.5 31.83$\pm$ 0.32 9.70$\pm$ 0.28 0.461$\pm$ 0.003 7.32$\pm$ 0.06 / 7.30$\pm$ 0.06 14.50$\pm$ 0.58 / 14.62$\pm$ 0.80 ODE 9.6 7.54$\pm$ 0.95 10.14$\pm$ 0.23 0.492$\pm$ 0.018 9.48$\pm$ 0.11 / 9.46$\pm$ 0.08 23.54$\pm$ 0.62 / 22.96$\pm$ 0.61 HYPRO 18.0 35.69$\pm$ x.xx 17.21$\pm$ x.xx 0.295$\pm$ x.xxx 6.63$\pm$ x.xx / 6.61$\pm$ x.xx 20.58$\pm$ x.xx / 20.53$\pm$ x.xx StackOverflow MostPopular 14.0 42.90$\pm$ 0.00 5.45$\pm$ 0.00 0.744$\pm$ 0.000 13.56$\pm$ 0.00 / 13.77$\pm$ 0.00 6.10$\pm$ 0.00 / 5.56$\pm$ 0.00 Last 10 9.3 26.42$\pm$ 0.00 5.20$\pm$ 0.00 0.934$\pm$ 0.000 14.52$\pm$ 0.00 / 14.55$\pm$ 0.00 8.67$\pm$ 0.00 / 6.72$\pm$ 0.00 IFTPP 24.1 45.41$\pm$ 0.11 13.00$\pm$ 0.79 0.641$\pm$ 0.002 13.57$\pm$ 0.07 / 13.64$\pm$ 0.05 8.78$\pm$ 0.87 / 8.31$\pm$ 0.50 IFTPP-K 15.3 44.85$\pm$ 0.24 11.16$\pm$ 1.07 0.644$\pm$ 0.003 13.41$\pm$ 0.07 / 13.51$\pm$ 0.06 12.42$\pm$ 0.97 / 11.42$\pm$ 0.78 RMTPP 14.5 45.43$\pm$ 0.13 13.33$\pm$ 0.25 0.701$\pm$ 0.007 12.95$\pm$ 0.02 / 13.17$\pm$ 0.05 13.26$\pm$ 0.29 / 12.72$\pm$ 0.16 RMTPP-K 12.0 44.89$\pm$ 0.09 11.72$\pm$ 0.10 0.689$\pm$ 0.001 12.92$\pm$ 0.01 / 13.13$\pm$ 0.01 14.91$\pm$ 0.19 / 14.30$\pm$ 0.09 NHP 13.0 44.53$\pm$ 0.05 10.86$\pm$ 0.19 0.715$\pm$ 0.004 13.02$\pm$ 0.02 / 13.24$\pm$ 0.02 12.67$\pm$ 0.55 / 11.96$\pm$ 0.40 AttNHP 15.5 45.17$\pm$ 0.10 12.67$\pm$ 0.13 0.705$\pm$ 0.001 13.08$\pm$ 0.03 / 13.30$\pm$ 0.02 11.95$\pm$ 0.19 / 11.13$\pm$ 0.32 ODE 13.9 44.38$\pm$ 0.09 10.12$\pm$ 0.19 0.711$\pm$ 0.004 13.04$\pm$ 0.02 / 13.27$\pm$ 0.03 11.37$\pm$ 0.32 / 10.52$\pm$ 0.23 HYPRO 11.8 45.18$\pm$ x.xx 12.88$\pm$ x.xx 0.715$\pm$ x.xxx 13.04$\pm$ x.xx / 13.26$\pm$ x.xx 15.57$\pm$ x.xx / 14.69$\pm$ x.xx
6 Experiments
We conduct experiments on five datasets of varying sizes and origins: the Transactions dataset (AI-Academy, 2021), the MIMIC-IV medical dataset (Johnson et al., 2023), two social networks datasets, Retweet (Zhao et al., 2015) and StackOverflow (Jure, 2014), and the Amazon reviews dataset (Jianmo, 2018). Dataset statistics and evaluation parameters are provided in Appendix C.4.
We evaluate representative modeling methods from different groups:
  1. 1.
    Rule-based baselines. MostPopular Generates a constant output with the most popular label from the prefix and the average time step. The Last N baseline outputs N previous events with adjusted timestamps.
  2. 2.

### v3

@@ 4. HoTPP Benchmark
Method Mean length Next-event Long-horizon Acc (%) mAP (%) MAE OTD Val / Test T-mAP Val / Test (%) Transactions MostPopular 5.0 32.69$\pm$ 0.00 1.09$\pm$ 0.00 0.649$\pm$ 0.000 7.03$\pm$ 0.00 / 7.05$\pm$ 0.00 2.86$\pm$ 0.00 / 3.10$\pm$ 0.00 HistoryDensity 154.2 32.78$\pm$ 0.00 0.85$\pm$ 0.00 1.063$\pm$ 0.000 7.42$\pm$ 0.00 / 7.44$\pm$ 0.00 7.92$\pm$ 0.00 / 7.51$\pm$ 0.00 IFTPP 13.3 38.27$\pm$ 0.03 4.65$\pm$ 0.03 0.637$\pm$ 0.001 6.79$\pm$ 0.01 / 6.81$\pm$ 0.01 6.20$\pm$ 0.10 / 6.08$\pm$ 0.14 IFTPP-T 12.8 37.68$\pm$ 0.03 4.30$\pm$ 0.03 0.640$\pm$ 0.001 6.84$\pm$ 0.02 / 6.85$\pm$ 0.02 5.80$\pm$ 0.27 / 5.55$\pm$ 0.19 RMTPP 8.7 38.28$\pm$ 0.04 4.64$\pm$ 0.02 0.693$\pm$ 0.005 6.68$\pm$ 0.01 / 6.70$\pm$ 0.01 7.41$\pm$ 0.35 / 7.12$\pm$ 0.10 NHP 9.4 35.47$\pm$ 0.12 3.36$\pm$ 0.02 0.693$\pm$ 0.005 6.99$\pm$ 0.03 / 6.99$\pm$ 0.03 5.66$\pm$ 0.17 / 5.48$\pm$ 0.12 AttNHP 6.8 30.68$\pm$ x.xx 1.09$\pm$ x.xx 0.749$\pm$ x.xxx 7.47$\pm$ x.xx / 7.48$\pm$ x.xx 1.50$\pm$ x.xx / 1.49$\pm$ x.xx ODE 9.1 35.60$\pm$ 0.06 3.34$\pm$ 0.06 0.695$\pm$ 0.002 6.96$\pm$ 0.01 / 6.97$\pm$ 0.01 5.53$\pm$ 0.08 / 5.52$\pm$ 0.13 HYPRO 7.3 22.50$\pm$ x.xx 4.62$\pm$ x.xx 0.694$\pm$ x.xxx 7.62$\pm$ x.xx / 7.62$\pm$ x.xx 8.20$\pm$ x.xx / 7.54$\pm$ x.xx Diffusion 10.4 36.05$\pm$ 0.07 3.66$\pm$ 0.03 0.650$\pm$ 0.002 6.86$\pm$ 0.01 / 6.88$\pm$ 0.00 5.95$\pm$ 0.05 / 6.04$\pm$ 0.08 MIMIC-IV MostPopular 2.7 4.80$\pm$ 0.00 2.75$\pm$ 0.00 14.52$\pm$ 0.00 19.75$\pm$ 0.00 / 19.75$\pm$ 0.00 1.63$\pm$ 0.00 / 1.56$\pm$ 0.00 HistoryDensity 160.0 0.74$\pm$ 0.00 2.63$\pm$ 0.00 3.23$\pm$ 0.00 18.94$\pm$ 0.00 / 18.92$\pm$ 0.00 4.25$\pm$ 0.00 / 4.19$\pm$ 0.00 IFTPP 12.4 58.52$\pm$ 0.06 46.04$\pm$ 0.26 3.00$\pm$ 0.02 11.53$\pm$ 0.03 / 11.55$\pm$ 0.03 22.56$\pm$ 0.13 / 22.31$\pm$ 0.15 IFTPP-T 11.7 58.57$\pm$ 0.04 47.96$\pm$ 1.51 3.03$\pm$ 0.01 11.46$\pm$ 0.02 / 11.48$\pm$ 0.02 24.07$\pm$ 0.31 / 23.68$\pm$ 0.29 RMTPP 9.8 58.24$\pm$ 0.14 45.62$\pm$ 1.28 3.93$\pm$ 0.03 13.69$\pm$ 0.05 / 13.76$\pm$ 0.06 22.00$\pm$ 0.28 / 21.62$\pm$ 0.29 NHP 3.0 23.19$\pm$ 0.61 11.67$\pm$ 0.39 4.71$\pm$ 0.31 17.92$\pm$ 0.34 / 17.92$\pm$ 0.33 9.76$\pm$ 0.41 / 9.80$\pm$ 0.38 AttNHP 6.2 44.99$\pm$ 1.11 31.88$\pm$ 0.64 2.95$\pm$ 0.20 14.30$\pm$ 0.16 / 14.31$\pm$ 0.16 19.98$\pm$ 0.79 / 19.74$\pm$ 0.71 ODE 11.6 44.16$\pm$ 0.76 27.48$\pm$ 1.04 2.91$\pm$ 0.01 14.36$\pm$ 0.20 / 14.39$\pm$ 0.20 16.37$\pm$ 0.75 / 16.19$\pm$ 0.74 HYPRO 7.4 45.69$\pm$ x.xx 45.54$\pm$ x.xx 3.98$\pm$ x.xx 14.68$\pm$ x.xx / 14.75$\pm$ x.xx 16.41$\pm$ x.xx / 16.13$\pm$ x.xx Diffusion 15.4 46.22$\pm$ 5.31 33.67$\pm$ 3.66 2.99$\pm$ 0.01 13.25$\pm$ 0.15 / 13.28$\pm$ 0.14 22.90$\pm$ 0.10 / 22.82$\pm$ 0.12 Retweet MostPopular 9.9 58.50$\pm$ 0.00 39.85$\pm$ 0.00 18.82$\pm$ 0.00 141.7$\pm$ 0.0 / 139.1$\pm$ 0.0 29.72$\pm$ 0.00 / 30.02$\pm$ 0.00 HistoryDensity 14.9 58.62$\pm$ 0.00 39.86$\pm$ 0.00 20.68$\pm$ 0.00 178.5$\pm$ 0.0 / 177.0$\pm$ 0.0 44.48$\pm$ 0.00 / 44.60$\pm$ 0.00 IFTPP 23.1 59.75$\pm$ 0.25 46.23$\pm$ 0.31 18.27$\pm$ 0.05 165.8$\pm$ 0.2 / 164.8$\pm$ 0.3 43.93$\pm$ 0.69 / 40.59$\pm$ 0.84 IFTPP-T 23.2 59.88$\pm$ 0.06 46.07$\pm$ 0.11 18.47$\pm$ 0.06 167.4$\pm$ 1.3 / 166.1$\pm$ 1.1 43.73$\pm$ 1.88 / 40.05$\pm$ 1.73 RMTPP 16.6 60.01$\pm$ 0.06 46.81$\pm$ 0.03 18.60$\pm$ 0.04 169.2$\pm$ 1.0 / 168.0$\pm$ 1.0 49.03$\pm$ 0.15 / 45.60$\pm$ 0.28 NHP 18.1 60.05$\pm$ 0.08 46.84$\pm$ 0.13 18.39$\pm$ 0.12 167.0$\pm$ 1.7 / 165.7$\pm$ 1.7 49.25$\pm$ 0.37 / 45.64$\pm$ 0.26 AttNHP 29.8 60.06$\pm$ 0.03 46.66$\pm$ 0.02 18.32$\pm$ 0.02 174.7$\pm$ 0.7 / 172.9$\pm$ 0.7 26.74$\pm$ 1.16 / 24.24$\pm$ 0.94 ODE 17.7 59.93$\pm$ 0.06 46.36$\pm$ 0.30 18.53$\pm$ 0.09 168.2$\pm$ 1.1 / 166.7$\pm$ 1.0 48.95$\pm$ 0.30 / 45.45$\pm$ 0.38 HYPRO 16.2 50.84$\pm$ x.xx 46.84$\pm$ x.xx 18.61$\pm$ x.xx 161.8$\pm$ x.x / 160.4$\pm$ x.x 50.58$\pm$ x.xx / 46.87$\pm$ x.xx Diffusion 14.7 57.72$\pm$ 0.77 42.96$\pm$ 0.67 18.60$\pm$ 0.06 159.0$\pm$ 1.4 / 158.0$\pm$ 1.1 54.24$\pm$ 0.89 / 52.24$\pm$ 0.68 Amazon MostPopular 5.0 33.46$\pm$ 0.00 9.58$\pm$ 0.00 0.304$\pm$ 0.000 6.33$\pm$ 0.00 / 6.30$\pm$ 0.00 10.10$\pm$ 0.00 / 9.72$\pm$ 0.00 HistoryDensity 71.8 32.68$\pm$ 0.00 9.63$\pm$ 0.00 0.325$\pm$ 0.000 7.29$\pm$ 0.00 / 7.27$\pm$ 0.00 28.25$\pm$ 0.00 / 27.53$\pm$ 0.00 IFTPP 12.8 35.54$\pm$ 0.08 16.72$\pm$ 0.04 0.250$\pm$ 0.001 6.70$\pm$ 0.01 / 6.64$\pm$ 0.01 23.05$\pm$ 0.48 / 23.79$\pm$ 0.39 IFTPP-T 13.1 35.27$\pm$ 0.04 16.46$\pm$ 0.07 0.245$\pm$ 0.001 6.64$\pm$ 0.02 / 6.59$\pm$ 0.02 22.81$\pm$ 0.10 / 23.42$\pm$ 0.13 RMTPP 16.4 35.85$\pm$ 0.03 17.21$\pm$ 0.04 0.292$\pm$ 0.002 6.67$\pm$ 0.03 / 6.62$\pm$ 0.03 19.65$\pm$ 0.23 / 19.99$\pm$ 0.18 NHP 9.9 16.58$\pm$ 3.26 11.19$\pm$ 0.18 0.438$\pm$ 0.009 8.54$\pm$ 0.28 / 8.52$\pm$ 0.30 27.19$\pm$ 0.35 / 26.95$\pm$ 0.37 AttNHP 12.4 22.21$\pm$ 5.64 8.22$\pm$ 0.39 0.526$\pm$ 0.075 7.63$\pm$ 0.12 / 7.61$\pm$ 0.14 19.29$\pm$ 1.45 / 18.59$\pm$ 1.52 ODE 9.4 10.44$\pm$ 2.84 10.28$\pm$ 0.57 0.473$\pm$ 0.031 9.23$\pm$ 0.17 / 9.23$\pm$ 0.18 23.73$\pm$ 0.36 / 23.13$\pm$ 0.46 HYPRO 17.8 21.47$\pm$ x.xx 17.14$\pm$ x.xx 0.288$\pm$ x.xxx 6.95$\pm$ x.xx / 6.95$\pm$ x.xx 20.47$\pm$ x.xx / 20.28$\pm$ x.xx Diffusion 13.3 30.57$\pm$ 1.28 13.36$\pm$ 0.49 0.247$\pm$ 0.000 6.56$\pm$ 0.03 / 6.52$\pm$ 0.04 30.08$\pm$ 0.25 / 30.29$\pm$ 0.32 StackOverflow MostPopular 9.8 42.90$\pm$ 0.00 5.45$\pm$ 0.00 0.602$\pm$ 0.000 13.12$\pm$ 0.00 / 13.15$\pm$ 0.00 8.97$\pm$ 0.00 / 8.16$\pm$ 0.00 HistoryDensity 74.0 42.51$\pm$ 0.00 5.61$\pm$ 0.00 0.669$\pm$ 0.000 17.30$\pm$ 0.00 / 17.29$\pm$ 0.00 16.23$\pm$ 0.00 / 14.90$\pm$ 0.00 IFTPP 24.1 45.31$\pm$ 0.10 11.59$\pm$ 0.70 0.646$\pm$ 0.003 13.56$\pm$ 0.19 / 13.66$\pm$ 0.19 9.68$\pm$ 0.55 / 8.87$\pm$ 0.59 IFTPP-T 23.0 45.40$\pm$ 0.04 12.78$\pm$ 0.29 0.639$\pm$ 0.001 13.51$\pm$ 0.07 / 13.61$\pm$ 0.05 9.62$\pm$ 0.15 / 8.81$\pm$ 0.20 RMTPP 12.7 45.30$\pm$ 0.07 12.61$\pm$ 0.40 0.717$\pm$ 0.006 13.00$\pm$ 0.03 / 13.25$\pm$ 0.04 13.81$\pm$ 0.30 / 13.35$\pm$ 0.28 NHP 11.9 44.12$\pm$ 0.79 9.90$\pm$ 1.16 0.735$\pm$ 0.021 13.27$\pm$ 0.31 / 13.46$\pm$ 0.27 13.12$\pm$ 0.26 / 12.41$\pm$ 0.31 AttNHP 15.6 45.17$\pm$ 0.16 12.56$\pm$ 0.15 0.703$\pm$ 0.000 13.10$\pm$ 0.03 / 13.31$\pm$ 0.01 11.95$\pm$ 0.38 / 11.15$\pm$ 0.29 ODE 13.3 44.19$\pm$ 0.13 9.08$\pm$ 0.27 0.728$\pm$ 0.007 13.12$\pm$ 0.05 / 13.35$\pm$ 0.05 11.48$\pm$ 0.29 / 10.75$\pm$ 0.21 HYPRO 11.8 30.54$\pm$ 0.70 12.61$\pm$ 0.40 0.718$\pm$ 0.006 14.00$\pm$ 0.19 / 14.12$\pm$ 0.17 15.25$\pm$ 0.28 / 14.84$\pm$ 0.23 Diffusion 14.4 37.03$\pm$ 1.24 7.48$\pm$ 0.16 0.668$\pm$ 0.006 12.87$\pm$ 0.17 / 13.01$\pm$ 0.16 15.80$\pm$ 0.85 / 15.07$\pm$ 0.74
Computational Efficiency. Some methods are particularly slow, especially during autoregressive inference on large datasets. Straightforward evaluation can take several hours on a single Nvidia V100 GPU. We optimized the training and inference pipelines in two ways to accelerate computation. First, we implemented an efficient RNN that reuses computations during parallel autoregressive inference from multiple starting positions. Second, we developed highly optimized versions of the thinning algorithm used for sampling in NHP and continuous-time neural architectures (NHP and ODE) (Rizoiu et al., 2017), achieving up to a 4x performance improvement compared to the official implementations. These optimizations allowed us to conduct the first large-scale evaluations of algorithms such as NHP, AttNHP, ODE, and HYPRO on datasets like Transactions and MIMIC-IV. Additionally, we provide the first open-source CUDA implementation of the batched linear assignment solver, significantly enhancing the applicability of the proposed T-mAP metric.
A detailed description of the benchmark can be found in Appendix C, with further details on HoTPP’s performance optimizations outlined in Appendix C.5.
4.1. Datasets
We for the first time combine datasets from a wide range of domains such as financial transactions, social networks, and medical records into a single evaluation benchmark. Specifically, we provide evaluation results on a transactional dataset (Babaev et al., 2022), the MIMIC-IV medical dataset (Johnson et al., 2023), and social network datasets (Retweet (Zhao et al., 2015), StackOverflow (Jure, 2014), and Amazon (Jianmo, 2018)). These datasets represent diverse underlying processes: social network data is influenced by cascades (Zhao et al., 2015), medical records exhibit repetitive patterns and transactional data reflects daily activities, combining regularity with significant uncertainty.
The dataset statistics are presented in Table 1. The Transactions dataset has the longest average sequence length and the largest number of classes, while MIMIC-IV contains the highest number of sequences. Retweet is medium size, whereas Amazon and StackOverflow can be considered small datasets. Further datasets details are reported in Appendix C.4 and Appendix F.
4.2. Methods
We implement representatives from different groups, listed below.
Statistical baselines. MostPopular generates K future events with average time steps. The predicted labels approximate distribution of labels in historical data. The HistoryDistribution baseline resembles MostPopular, but estimates types densities w.r.t. time features.